# bibliography.bib

### v1

@@ Abstract.
Thanks to Dan Appel, Reid Dale, Noah Schweber, Ted Slaman, and Kentarô Yamamoto for helpful discussion. Thanks to Neil Barton, Wes Holliday, and Andrew Lee for comments on drafts.
@@ 1. Introduction
This completeness theorem falls out of an exact characterization of the sets (and, in general, the $n$-ary relations) that are modally definable over $\mathbf{FIX}$. From this characterization, we infer that the set of intrinsic sentences is not modally indefinable. Thus, it is not a mere coincidence that Kripke does not lapse into modal language when defining intrinsicality. Indeed, we are able to show that the intrinsic sentences are indiscernible within the modal language from another set, namely, the set of inevitable sentences, which we define later in the paper. The characterization of the modally definable sets builds on a classic normal form theorem for $\mathbf{S5}$ from [carnap1946modalities]. According to Carnap’s theorem, every formula is $\mathbf{S5}$-equivalent to a disjunction of formulas in a very special form. This normal form result and its corollaries adumbrated here resemble classic results in model theory (particular, results in o-minimality and quantifier elimination). We conclude with a discussion of this analogy.
@@ 5.2. Matrices
Consider Table 3, Table 4, Table 5, and Table 6. Each of these images exhibits a subset of the $3\times 3$ matrix. In each case, consider the following question: Is there a variable assignment $\star$ such that $x_{1}^{\star}$ and $x_{2}^{\star}$ can attain all and only these combinations of values? Put another way, given a subset $\mathcal{S}$ of the $3\times 3$ matrix, is there a variable assignment $\star$ such that $\mathbf{FIX}_{V^{\star}}\Vdash\bigwedge\{\lozenge\varphi\mid\varphi\in\mathcal{S}\}\wedge\bigwedge\{\neg\lozenge\varphi\mid\varphi\notin\mathcal{S}\}$?
@@ 5.4. The Main Lemma
We prove the contrapositive. Suppose $\mathbf{FIX}\Vdash_{V^{\star}}\lozenge\varphi$. Then there is some $w$ such that $\mathbf{FIX},w\Vdash_{V^{\star}}\varphi$. We now break into cases depending on whether any of the $C_{s}$ is true at $w$.
@@
Subsubcase 2.2.1: Suppose $s\in\mathcal{S}^{\Phi}$. Then $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\leftrightarrow A^{i}_{s}$, whence $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\leftrightarrow\theta$. We reason as follows:
$\varphi_{i}\text{ is }T(x_{i})$ $\Longleftrightarrow\mathbf{FIX},w\Vdash_{V^{\star}}T(x_{i})$
$\Longleftrightarrow\mathbb{N},w\Vdash_{\mathsf{K3}}\mathsf{True}(\ulcorner A^{i}\urcorner)$
$\Longleftrightarrow\mathbb{N},w\Vdash_{\mathsf{K3}}A^{i}$
$\Longleftrightarrow\mathbb{N},w\Vdash_{\mathsf{K3}}\theta$
$\Longleftrightarrow\theta\text{ is }0=0$
$\Longleftrightarrow\Phi(s)_{i}=T(x_{i})$
@@ 5.5. Completeness
$\mathbf{S5[Con,Ground,Min_{n}]}\vdash\neg\varphi\leftrightarrow(\psi_{1}\vee\dots\psi_{k})$
where each $\psi_{i}$ is a disjunction of $n$-isolators. Since $\psi$ is consistent, at least one of the disjuncts $\psi_{i}$ is also consistent. It follows that $\psi_{i}$ meets the prime conditions. By Lemma 5.9, there is an $\mathcal{L}_{T}$ realization $\star$ such that
@@ 5.6. Counting Definable Relations
How many $\mathbf{FIX}$-definable $n$-ary relations are there? One could presumably concoct a formula using the inclusion–exclusion principle. Yet it would probably be difficult to write down an informative one. The following coarse estimate proposition follows immediately:
@@ 6. Conclusions
The decidable/undecidable dichotomy is not exactly the same as the model-theoretic tame/wild dichotomy, but in these particular instances they coincide. In particular, Tarski’s theorem follows from o-minimality. A structure is o-minimal if every definable set is a finite union of intervals and a theory is o-minimal if all its structures are. Of course, there is much that o-minimal theories cannot express. The first-order theory of the reals, for instance, cannot define the set of integers or define the sine function. However, this permits logicians to give sharp, detailed characterizations of definable sets.

### v2

@@ Abstract.
Thanks to Dan Appel, Reid Dale, Noah Schweber, Ted Slaman, and Kentarô Yamamoto for helpful discussion. Thanks to Neil Barton, Wes Holliday, Andrew Lee, and Matt Mandelkern for comments on drafts.
@@ 1. Introduction
This completeness theorem falls out of an exact characterization of the sets (and, in general, the $n$-ary relations) that are modally definable over $\mathbf{FIX}$. From this characterization, we infer that the set of intrinsic sentences is not modally definable. Thus, it is not a mere coincidence that Kripke does not lapse into modal language when defining intrinsicality. Indeed, we are able to show that the intrinsic sentences are indiscernible within the modal language from another set, namely, the set of inevitable sentences, which we define later in the paper. The characterization of the modally definable sets builds on a classic normal form theorem for $\mathbf{S5}$ from [carnap1946modalities]. According to Carnap’s theorem, every formula is $\mathbf{S5}$-equivalent to a disjunction of formulas in a very special form. This normal form result and its corollaries adumbrated here resemble classic results in model theory (particular, results in o-minimality and quantifier elimination). We conclude with a discussion of this analogy.
@@ 5.2. Matrices
Consider Table 3, Table 4, Table 5, and Table 6. Each of these images exhibits a subset of the $3\times 3$ matrix. In each case, consider the following question: Is there a variable assignment $\star$ such that $x_{1}^{\star}$ and $x_{2}^{\star}$ can attain all and only these combinations of values? Put another way, given a subset $\mathcal{S}$ of the $3\times 3$ matrix, is there a variable assignment $\star$ such that $\mathbf{FIX}\Vdash_{V^{\star}}\bigwedge\{\lozenge\varphi\mid\varphi\in\mathcal{S}\}\wedge\bigwedge\{\neg\lozenge\varphi\mid\varphi\notin\mathcal{S}\}$?
@@ 5.4. The Main Lemma
We prove the contrapositive. Suppose $\mathbf{FIX}\Vdash_{V^{\star}}\lozenge\varphi$. Then there is some $w$ such that
$(\triangle)\quad\mathbf{FIX},w\Vdash_{V^{\star}}\varphi.$
We now break into cases depending on whether any of the $C_{s}$ is true at $w$.
@@
Subsubcase 2.2.1: Suppose $s\in\mathcal{S}^{\Phi}$. Then $\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\leftrightarrow A^{i}_{s}$, where $A^{i}_{s}$ is $C_{s}\to\theta$.
Note that, by the definition of $A^{i}_{s}$:
$(\circledast)\quad\theta\text{ is }0=0\text{ just in case }\Phi(s)_{i}=T(x_{i}).$
Since $\mathbb{N},w\vDash_{\mathsf{K3}}C_{s}$, we infer that:
$(\boxminus)\quad\mathbb{N},w\vDash_{\mathsf{K3}}A^{i}\leftrightarrow\theta.$
We reason as follows:
$\varphi_{i}\text{ is }T(x_{i})$ $\Longleftrightarrow\mathbf{FIX},w\Vdash_{V^{\star}}T(x_{i})\text{by ($\triangle)$}$
$\Longleftrightarrow\mathbb{N},w\Vdash_{\mathsf{K3}}A^{i}\text{ since $x_{i}^{\star}=A^{i}$}$
$\Longleftrightarrow\mathbb{N},w\Vdash_{\mathsf{K3}}\theta\text{ by ($\boxminus$)}$
$\Longleftrightarrow\theta\text{ is }0=0\text{ since $\theta\in\{0=0,0=1,\lambda\}$}$
$\Longleftrightarrow\Phi(s)_{i}=T(x_{i})\text{ by ($\circledast$)}$
@@ 5.5. Completeness
$\mathbf{S5[Con,Ground,Min_{n}]}\vdash\neg\varphi\leftrightarrow(\psi_{1}\vee\dots\vee\psi_{k})$
where each $\psi_{i}$ is a disjunction of $n$-isolators. Since $\neg\varphi$ is consistent, at least one of the disjuncts $\psi_{i}$ is also consistent. It follows that $\psi_{i}$ meets the prime conditions. By Lemma 5.9, there is an $\mathcal{L}_{T}$ realization $\star$ such that
@@ 5.6. Counting Definable Relations
How many $\mathbf{FIX}$-definable $n$-ary relations are there? One could presumably concoct a formula using the inclusion–exclusion principle. Yet it would probably be difficult to write down an informative one. The following coarse estimate follows immediately:
@@ 6. Conclusions
The decidable/undecidable dichotomy is not exactly the same as the model-theoretic tame/wild dichotomy, but in these particular instances they coincide. In particular, Tarski’s theorem follows from o-minimality. A structure is o-minimal if every definable set is a finite union of intervals333‘Interval’ is understood in a broad sense to include rays and points. and a theory is o-minimal if all its structures are. Of course, there is much that cannot be expressed over o-minimal structures. The first-order language of rings, for instance, cannot define the set of integers or the sine function over the reals. However, this permits logicians to give sharp, detailed characterizations of definable sets.